OpenStudy (anonymous):

If the equation y=2(2.1)^x is plotted on a semilog scale graph, then the linearized equation will be log(y)=mx+b where m=? b=?

OpenStudy (lalaly):

\[\log(y)=\log2(2.1)^x\]

OpenStudy (lalaly):

\[logy=\log2+\log2.1^x\]

OpenStudy (lalaly):

\[\log(y)=xlog(2.1) + \log2\]

OpenStudy (lalaly):

m=log2.1 b=log2